Construct a table of values to investigate the following limits.


1. lim (10/x-2)

x-> 7


2. lim (x+6/x+2)

x-> 1


3. lim f(x)

x-> -1

If f(x) { 1/x if x ≤ -1}

If f(x) { x2-2 if x > -1}



1
Expert's answer
2022-03-02T11:48:41-0500

1. The following table gives values of for values of close to "7," but not

equal to "7."


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & f(x) \\ \hline
 6 & 2.500000 \\
 \hdashline
 6.5 & 2.222222 \\
 \hdashline
 6.9 & 2.040816\\
 \hdashline
 6.99 & 2.004008 \\
 \hdashline
 6.999 & 2.000400 \\
 \hdashline
 6.9999 & 2.000040 \\
 \hdashline
 7.0001 & 1.999960 \\
 \hdashline
 7.001 & 1.999600 \\
 \hdashline
 7.01 & 1.996008 \\
 \hdashline
 7.1 & 1.960784 \\
 \hdashline
 7.5 & 1.818182 \\
 \hdashline
 8 & 1.666667 \\
 \hdashline
\end{array}"

"\lim\limits_{x\to 7}\dfrac{10}{x-2}=2"



2. The following table gives values of for values of close to "2," but not

equal to "1."


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & f(x) \\ \hline
 0 & 3.000000 \\
 \hdashline
 0.5 & 2.600000 \\
 \hdashline
 0.9 & 2.379310\\
 \hdashline
 0.99 & 2.337793 \\
 \hdashline
 0.999 & 2.333778 \\
 \hdashline
 0.9999 & 2.333378 \\
 \hdashline
 1.0001 & 2.333289 \\
 \hdashline
 1.001 & 2.332889 \\
 \hdashline
 1.01 & 2.328904 \\
 \hdashline
 1.1 & 2.290323 \\
 \hdashline
 1.5 & 2.142857 \\
 \hdashline
 2 & 2.000000 \\
 \hdashline
\end{array}"

"\lim\limits_{x\to 1}\dfrac{x+6}{x+2}=\dfrac{7}{3}"

3. The following table gives values of for values of close to "-1," but not

equal to "-1."


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & f(x) \\ \hline
 -2 & -0.500000 \\
 \hdashline
 -1.5 & -0.666667 \\
 \hdashline
 -1.1 & -0.909091\\
 \hdashline
 -1.01.99 & -0.990099 \\
 \hdashline
 -1.001 & -0.999001\\
 \hdashline
 -1.0001 & -0.999900 \\
 \hdashline
 -0.9999 & -1.000200 \\
 \hdashline
 -0.999 & -1.001999 \\
 \hdashline
 -0.99 & -1.0199\\
 \hdashline
 -0.9 & -1.19 \\
 \hdashline
 -0.5 & -1.75 \\
 \hdashline
 0 & -2 \\
 \hdashline
\end{array}"

"\lim\limits_{x\to -1}f(x)=-1"